OpenStudy (anonymous):

If f(x) = (x^2+1)^x, then f'(x) = ?

OpenStudy (solomonzelman):

It's asking you to derive (x^2+1)^x

OpenStudy (anonymous):

can someone post the steps to solving this problem?

OpenStudy (solomonzelman):

http://www.derivative-calculator.net/

OpenStudy (solomonzelman):

paste your problem there....

OpenStudy (anonymous):

thanks you

OpenStudy (zarkon):

Pretty soon we will not have to think at all ;)

OpenStudy (anonymous):

(d/dx) (x^n) = (n)(x^(n-1)) very handy. (d/dx)(x^3) = 3x^2.

OpenStudy (anonymous):

\[f(x)=(x^2+1)^x\] Logarithmic differentiation: \[\ln f(x)=\ln (x^2+1)^x\\ \ln f(x)=x\ln (x^2+1)\] Differentiate both sides with respect to \(x\): \[\frac{f'(x)}{f(x)}=\ln(x^2+1)+x\left(\frac{2x}{x^2+1}\right)\] Solve for \(f'(x)\).